ICM 1998, Berlin, Aug. 18–27

Abstracts of Short Communications and Posters

Section:

9. Ordinary Differential Equations and Dynamical Systems

1991 MS Classification: not given Ferreira, Fernando Flavio, Universidade Moderna Porto, Portugal, Pinto, A. A., Rand, D. A. C^{1+HD} bounds on the smoothness of Denjoy maps The smoothness of the holonomy for smooth attractors is bounded above by one plus the Hausdorff dimension of the invariant set along the stable leaves. The smoothness of denjoy maps which are fixed points of renormalization is bounded by one plus the Hausdorff dimension of the non-wandering set.